# J-HERTz: J-PLUS Heritage Exploration of Radio Targets at z 5

### v1

@@ 2.5 Object classification
We use the recently developed Bayesian Artificial Neural Networks for the Javalambre Observatory Survey (BANNJOS) machine learning-based algorithm (delPino2024) to classify sources into galaxies, QSOs, and stars. This algorithm uses Bayesian Artificial Neural Networks to provide a PDF (Probability Distribution Function) and correlations between the three classes. It uses the full photometric information provided by J-PLUS in different apertures, as well as the morphology and other external properties such as the proper motions from Gaia and the IR bands from WISE. Note that this classification refers to the optical/IR morphological and spectral properties of the counterparts and is not intended to classify the radio properties of the active nucleus. This method improves upon previous works because it includes uncertainties and correlations, allowing for a more precise filtering of the results according to the user’s preferences for accuracy and completeness.
@@
2.6 Radio emission sizes and fluxes
@@
When the source has a clear, continuous radio emission, Hardcastle2023 estimate the angular size with a Gaussian fit, using the resolution criterion described (Shimwell2022), where they use a integrated-to-peak flux ratio and a signal-to-noise threshold. When the source has several composite radio components they try to estimate the angular size from the largest dimension of the convex perimeter encompassing all components. A small number of sources also have manual size measurements, and others have a “flood-fill size” measurement, where they use the individual components as the starting point and only include pixels above the local noise level. All sources whose size is not measured with a Gaussian fit are taken to be resolved. There will be a higher resolution release in the future that will allow to better constrain compact sources. Using the Resolved flag in their catalog, there are 450k unresolved and only 40k resolved radio sources in our sample, but their criteria is very restrictive and only considers the most ideal cases as resolved. We decide to include all sources in our catalog and our analysis, but we only consider as valid the size measurements of sources with Resolved = True, which we will refer to simply as Resolved radio sources from now on. We include this flag in our final catalog as well. LoTSS provides the peak flux density as well as the total flux within the estimated angular size of a source.
2.7 Redshifts
@@
There are available spectroscopic redshifts for 34% of our sources, mainly from SDSS and DESI, and photometric redshift estimates based on broad bands for 95% of sources in LoTSS and narrow bands for 97% of sources in J-PLUS. Hernan2024 demonstrate that photo-z accuracy can be improved when the probability distributions from a deeper wide-band survey are combined with those from a shallower narrow-band survey. This method outperforms alternatives such as fitting all the photometry together or using a weighted average of point estimates.
@@
LoTSS DR2 includes spectroscopic redshifts from SDSS DR16, the first data release of the DESI spectroscopic survey, or HETDEX DR1 when available. We update or add new SDSS spectroscopic redshifts from the latest DR18 performing a 2 arcsec cross-match between the SDSS positions and the J-PLUS optical/IR identifications in catWISE and SDSS. When the spectroscopic redshift from SDSS/DESI is above 5, or has a difference with the photometric redshift of that same survey larger than 0.2, it is marked with the unreliable_spec_z flag. LoTSS also provides photometric redshift estimates for the sample with optical identifications in the DESI Legacy Imaging Surveys DR8 taken from Duncan2022. The latter uses Gaussian Mixture Models (GMMs) to separate the parameter space into different regions that are trained and tested separately with the GPz code (Almosallam2016). They include a ‘best z’, which always contains the spectroscopic redshift if available, or the photo-z estimate if flagged as reliable. If none of these are available, it is left empty. The J-PLUS DR3 photometric system includes essential broad and narrow bands that enable precise photo-z measurements for nearby galaxies with (Molino2019). This accuracy is achieved through continuous coverage of the 4 000 Å break and the detection of specific emission lines. The photo-z estimates are obtained the same way as with the miniJPAS catalog (Bonoli2021), which is described in detail in Hernan2021, but adapted to the properties of the J-PLUS data. In short, they use a modified version of the LePhare code (Arnouts1999) and input 50 synthetic galaxy templates generated with the Code Investigating GALaxy Emission (CIGALE; Boquien2019). They search in the range of 0 z 1 with a resolution of 0.001 to prioritize physically plausible solutions, as high-redshift objects in J-PLUS are expected to be dominated by QSOs due to brightness limits. They apply a prior derived from galaxy counts in the VIMOS VLT Deep Survey (VVDS LeFevre2005) and obtain final redshift PDFs for all sources regardless of their morphological classification to avoid biases, albeit only galaxy templates are used. Therefore, the J-PLUS photo-zs should be primarily trusted only for sources classified as galaxies.
@@
Both J-PLUS and LoTSS provide PDFs for the photo-z estimates. In the latter, the GMM method assumes the PDF is Gaussian by definition, where the mean is the photo-z estimate and the standard deviation its error. In J-PLUS, the PDF can have an arbitrary shape and even be multi-modal. The photo-z estimate is the value with the maximum and its error is obtained from the 68th confidence interval. By leveraging the datasets from both surveys, we can enhance precision by conflating their PDFs to derive a new combined photo-z estimation, following Hernan2024. We only use the estimates of objects classified as galaxies and photo-z in both surveys, as the J-PLUS photo-z estimates are computed using lower redshift galaxy templates. Among these, the percentage of sources with a photo-z confidence interval below 0.05 increases from 33% in J-PLUS and 56% in LoTSS to 72% in the new combined photo-z. The number of combined photo-zs that have smaller errors than either survey is 235,000. In cases with available spectroscopic redshifts, Fig. 4 shows the comparison with the different photo-zs for galaxies with both photo and spectroscopic redshifts below . To quantify the improvement in the redshift determinations, the following statistical diagnostics are introduced:
@@
The latest SDSS DR18 (Almeida2023) spectroscopic redshift, if available, without warnings and unreliable_spec_z = 0.
@@
where J-PLUS ODDS is an indicator of the confidence in and is defined as the probability that is smaller than 0.03. LoTSS flag_qual selects sources with reliable redshifts, reasonable uncertainty, minimal contamination from nearby sources, low star-likelihood and free from imaging artifacts. We provide the z_source column, which specifies the origin of the redshift in each case, and can be one of the following: ‘Combined zphot’, ‘J-PLUS zphot’, ‘LoTSS zphot’, ‘LoTSS zspec’, and ‘SDSS DR18 zspec’. If none of the listed conditions are fulfilled, z_best and z_source are left blank. Roughly 34% of total sources in the catalog have a spectroscopic best redshift, and 92% of the galaxies have a best combined photo-z. We also provide an incompatible_spec_class flag, which indicates when the SDSS/DESI spectroscopic class is different from the BANNJOS class (more information in Section 4.4). Our z_best together with the cosmology described at the end of Section 1 is used to estimate the radio linear size from the estimated angular size and the radio luminosity () from the total source flux on the assumption of a spectral index = 0.7.
2.8 Radio-loudness
@@
Traditionally, AGN are divided into radio-loud and radio-quiet classes. The former correspond to the kinetic mode of black hole activity, where most of the energy output is dominated by powerful jets (Merloni2007). On the other hand, radio-quiet sources are in the radiative mode, dominated by the accretion disk, although the origin of their radio emission is not well established yet (Chen2023). There are various radio-loudness definitions in the literature, e.g., based on the radio-to-optical (Kellermann1989) or radio-to-X-ray (Terashima2003) ratios. Here we introduce a new radio-loudness definition, based on the flux ratio between 144 MHz and WISE W2 (), to analyze the distribution of star-forming galaxies and AGN in our sample. This ratio, discussed in this work for the first time, is particularly convenient for identifying radio-loud AGN, including optically quiescent radio galaxies with past jet activity, because the low-frequency radio flux is sensitive to older and less energetic synchrotron emission. On the other hand, the IR flux probes the presence of hot dust associated with the active nucleus or star formation. In star-forming galaxies this ratio would be dominated by the star-formation activity and the associated contribution from supernova remnants, while in radio-quiet QSOs low ratios are caused by the strong IR dust emission. Radio-loud sources have an extra contribution to the radio fluxes caused by the jet activity, which, at low frequencies typically comes from extended radio structures. This extra contribution distinguishes the galaxies with high radio-loudness ratios (hereafter radio-loud galaxies) from star-forming galaxies and radio-quiet QSOs. Note that although the rest-frame frequencies of this ratio are different for each redshift, our range is not large enough to avoid tracing the synchrotron continuum in radio and the dust emission in the IR. Most galaxies and a large fraction of QSOs are below . In this case, the WISE W2 band rest-frequency would correspond to , which is still representative of the AGN hot dust (e.g. Elvis1994). On the other hand, the LoTSS 144 MHz frequency would correspond to approximately 500 MHz, which still probes the synchrotron continuum. Although redshift may introduce a scatter in the radio-to-IR ratios, the latter still provides a meaningful radio-quiet and radio-loud source separation. Additionally, variability may affect the low frequency and mid- to near IR continuum differently, owing to the distinct timescales associated with the jet and hot dust components. A larger variability is expected in the latter due to the faster response to luminosity changes in the accretion disk. Nevertheless, we only expect a minor fraction of IR-variable sources in our sample (10%), as suggested by Sanchez2017 for a sample of AGN at z 2. Additionally, the radio-loudness distribution covers a range of 7 orders of magnitude, thus significant deviations are only expected for sources with extraordinarily large variability amplitudes.
2.9 Stellar masses and star formation rates
@@
Arizo2025 obtain stellar masses, SFRs, rest-frame colors and luminosities from J-PLUS extended sources according to the sglc_prob_star parameter (LopezSanJuan2019) with the CIGALE SED fitting code. They adopt J-PLUS photometric redshifts, fluxes and flux errors and use Bruzual2003 stellar populations with Chabrier2003 initial mass functions with different metallicities. We scale their mass estimates and SFRs (averaged over the last 100 Myrs), initially computed for sources with J-PLUS photo-z estimates up to , to our new z_best using the previously defined cosmology. Approximately % of the sources in our sample have a stellar mass estimate, with the majority (95%) being galaxies. We discard the values for QSOs, as AGN templates in CIGALE do not provide reliable stellar mass estimates.
@@
These results are compared with LoTSS-based estimates from Hardcastle2023, who follow the SED-fitting approach in Duncan2022 using DESI Legacy Surveys broad-bands and WISE photometry, but introducing photo-z uncertainties. After scaling their masses to our z_best, both estimates are compared in the left panel of Fig. 5. LoTSS values show a positive median offset of 0.18 dex with 0.20 dex scatter for galaxies with reliable estimates and . The offset likely arises from differences in flux measurements and apertures, while the scatter reflects photometric uncertainties. Brighter galaxies ( and z 0.2) show a lower scatter than the median squared error of the J-PLUS mass estimates, meaning that the errors are slightly overestimated. Both offset and scatter are slightly larger for fainter galaxies ( and z 0.2). Overall, J-PLUS narrow-band derived masses are in good agreement with LoTSS broad-band estimates, confirming the robustness of our approach. The J-HERTz catalog includes stellar masses and SFRs at z_best from J-PLUS photometry. The right panel of Fig. 5 shows the distribution of stellar masses, with a few extreme outliers caused by large differences between the original J-PLUS photo-z and the newly obtained z_best, which mainly happens whtn the spectroscopic class in SDSS/DESI is not a galaxy and can be filtered with the incompatible_spec_class flag.
3 Source properties in J-HERTz
3.1 QSOs
@@
In this section we analyze the main properties of the nearly 31k QSOs included in the catalog. These include their best redshift, radio luminosity, radio linear size and radio-loudness, whose distributions are shown in the orange histograms of Fig. 6. The QSO best-z distribution in our sample peaks around , in agreement with the higher density of active galaxies found at the cosmic noon (MadauDickinson2014), but it spans a wide range of values. The lowest value with a compatible spectroscopic class is 0.028 from SDSS (WISEA J160443.63+531517.8), although this appears to be a misidentification of H emission, which actually corresponds to [OII] emission at . The highest value of also comes from SDSS (WISEA J234219.09+270845.9), however spectroscopic redshifts above 5 are unreliable. Cases like either of these two examples (when the spec-z and photo-z of the same spectroscopic survey are very different, or the spec-z is larger than 5) are marked with an unreliable_spec_z flag and are not included in the histogram or the z_best parameter, although they are included in the zspec parameter (see Section 4.4). The radio luminosities peak at , with the majority falling between and . The few outliers with very low luminosities correspond to sources where the z_best comes from a low spec_z of an object spectrally classified as a star, and can be avoided with the incompatible_spec_class flag. The estimated linear size of the QSO radio emission for Resolved sources with an angular size larger than 6 arcsec mostly falls below 2700 kpc, with the most extended radio source reaching 4.3 Mpc and 20 arcmin (WISEA J093139.04+320400.1; Kuzmicz2018). We consider the host galaxies resolved if their major axis () is larger than the typical FWHM value of 1.2 arcsec in J-PLUS. By this criterion, approximately 20% of the QSO host galaxies detected by J-PLUS are resolved. Over 70% of these have an axis ratio , indicating a high degree of sphericity typical of elliptical galaxies, which are likely the usual hosts of QSOs. Taking only the cases with a relative error in and bands below 10%, QSOs have a median and dispersion () color of 0.20 0.12. The radio-loudness of QSOs is not described by a single Gaussian distribution, showing instead a bimodal distribution, i.e. with two local maxima centered around / and . We explore this further in Section 4.1.
3.2 Galaxies

### v2

@@ II.5 Object classification
We use the recently developed Bayesian Artificial Neural Networks for the Javalambre Observatory Survey (BANNJOS) machine learning-based algorithm (del Pino et al., 2024) to classify sources into galaxies, QSOs, and stars. This algorithm uses Bayesian Artificial Neural Networks to provide a PDF (Probability Distribution Function) and correlations between the three classes. It uses the full photometric information provided by J-PLUS in different apertures, as well as the morphology and other external properties such as the proper motions from Gaia and the IR bands from WISE. Note that this classification refers to the optical/IR morphological and spectral properties of the counterparts and is not intended to classify the radio properties of the active nucleus. This method improves upon previous works because it includes uncertainties and correlations, allowing for a more precise filtering of the results according to the user’s preferences for accuracy and completeness.
@@
II.6 Radio emission sizes and fluxes
@@
When the source has a clear, continuous radio emission, Hardcastle et al. (2023) estimate the angular size with a Gaussian fit, using the resolution criterion described (Shimwell et al., 2022), where they use a integrated-to-peak flux ratio and a signal-to-noise threshold. When the source has several composite radio components they try to estimate the angular size from the largest dimension of the convex perimeter encompassing all components. A small number of sources also have manual size measurements, and others have a “flood-fill size” measurement, where they use the individual components as the starting point and only include pixels above the local noise level. All sources whose size is not measured with a Gaussian fit are taken to be resolved. There will be a higher resolution release in the future that will allow to better constrain compact sources. Using the Resolved flag in their catalog, there are 450k unresolved and only 40k resolved radio sources in our sample, but their criteria is very restrictive and only considers the most ideal cases as resolved. We decide to include all sources in our catalog and our analysis, but we only consider as valid the size measurements of sources with Resolved = True, which we will refer to simply as Resolved radio sources from now on. We include this flag in our final catalog as well. LoTSS provides the peak flux density as well as the total flux within the estimated angular size of a source.
II.7 Redshifts
@@
There are available spectroscopic redshifts for 34% of our sources, mainly from SDSS and DESI, and photometric redshift estimates based on broad bands for 95% of sources in LoTSS and narrow bands for 97% of sources in J-PLUS. Hernán-Caballero et al. (2024) demonstrate that photo-z accuracy can be improved when the probability distributions from a deeper wide-band survey are combined with those from a shallower narrow-band survey. This method outperforms alternatives such as fitting all the photometry together or using a weighted average of point estimates.
@@
LoTSS DR2 includes spectroscopic redshifts from SDSS DR16, the first data release of the DESI spectroscopic survey, or HETDEX DR1 when available. We update or add new SDSS spectroscopic redshifts from the latest DR18 performing a 2 arcsec cross-match between the SDSS positions and the J-PLUS optical/IR identifications in catWISE and SDSS. When the spectroscopic redshift from SDSS/DESI is above 5, or has a difference with the photometric redshift of that same survey larger than 0.2, it is marked with the unreliable_spec_z flag. LoTSS also provides photometric redshift estimates for the sample with optical identifications in the DESI Legacy Imaging Surveys DR8 taken from Duncan (2022). The latter uses Gaussian Mixture Models (GMMs) to separate the parameter space into different regions that are trained and tested separately with the GPz code (Almosallam et al., 2016). They include a ‘best z’, which always contains the spectroscopic redshift if available, or the photo-z estimate if flagged as reliable. If none of these are available, it is left empty. The J-PLUS DR3 photometric system includes essential broad and narrow bands that enable precise photo-z measurements for nearby galaxies with (Molino et al., 2019). This accuracy is achieved through continuous coverage of the 4 000 Å break and the detection of specific emission lines. The photo-z estimates are obtained the same way as with the miniJPAS catalog (Bonoli et al., 2021), which is described in detail in Hernán-Caballero et al. (2021), but adapted to the properties of the J-PLUS data. In short, they use a modified version of the LePhare code (Arnouts et al., 1999) and input 50 synthetic galaxy templates generated with the Code Investigating GALaxy Emission (CIGALE; Boquien et al., 2019). They search in the range of 0 z 1 with a resolution of 0.001 to prioritize physically plausible solutions, as high-redshift objects in J-PLUS are expected to be dominated by QSOs due to brightness limits. They apply a prior derived from galaxy counts in the VIMOS VLT Deep Survey (VVDS Le Fèvre et al., 2005) and obtain final redshift PDFs for all sources regardless of their morphological classification to avoid biases, albeit only galaxy templates are used. Therefore, the J-PLUS photo-zs should be primarily trusted only for sources classified as galaxies.
@@
Both J-PLUS and LoTSS provide PDFs for the photo-z estimates. In the latter, the GMM method assumes the PDF is Gaussian by definition, where the mean is the photo-z estimate and the standard deviation its error. In J-PLUS, the PDF can have an arbitrary shape and even be multi-modal. The photo-z estimate is the value with the maximum and its error is obtained from the 68th confidence interval. By leveraging the datasets from both surveys, we can enhance precision by conflating their PDFs to derive a new combined photo-z estimation, following Hernán-Caballero et al. (2024). We only use the estimates of objects classified as galaxies and photo-z in both surveys, as the J-PLUS photo-z estimates are computed using lower redshift galaxy templates. Among these, the percentage of sources with a photo-z confidence interval below 0.05 increases from 33% in J-PLUS and 56% in LoTSS to 72% in the new combined photo-z. The number of combined photo-zs that have smaller errors than either survey is 235,000. In cases with available spectroscopic redshifts, Fig. 4 shows the comparison with the different photo-zs for galaxies with both photo and spectroscopic redshifts below . To quantify the improvement in the redshift determinations, the following statistical diagnostics are introduced:
@@
The latest SDSS DR18 (Almeida et al., 2023) spectroscopic redshift, if available, without warnings and unreliable_spec_z = 0.
@@
where J-PLUS ODDS is an indicator of the confidence in and is defined as the probability that is smaller than 0.03. LoTSS flag_qual selects sources with reliable redshifts, reasonable uncertainty, minimal contamination from nearby sources, low star-likelihood and free from imaging artifacts. We provide the z_source column, which specifies the origin of the redshift in each case, and can be one of the following: ‘Combined zphot’, ‘J-PLUS zphot’, ‘LoTSS zphot’, ‘LoTSS zspec’, and ‘SDSS DR18 zspec’. If none of the listed conditions are fulfilled, z_best and z_source are left blank. Roughly 34% of total sources in the catalog have a spectroscopic best redshift, and 92% of the galaxies have a best combined photo-z. We also provide an incompatible_spec_class flag, which indicates when the SDSS/DESI spectroscopic class is different from the BANNJOS class (more information in Section IV.4). Our z_best together with the cosmology described at the end of Section I is used to estimate the radio linear size from the estimated angular size and the radio luminosity () from the total source flux on the assumption of a spectral index = 0.7.
II.8 Radio-loudness
@@
Traditionally, AGN are divided into radio-loud and radio-quiet classes. The former correspond to the kinetic mode of black hole activity, where most of the energy output is dominated by powerful jets (Merloni and Heinz, 2007). On the other hand, radio-quiet sources are in the radiative mode, dominated by the accretion disk, although the origin of their radio emission is not well established yet (Chen et al., 2023). There are various radio-loudness definitions in the literature, e.g., based on the radio-to-optical (Kellermann et al., 1989) or radio-to-X-ray (Terashima and Wilson, 2003) ratios. Here we introduce a new radio-loudness definition, based on the flux ratio between 144 MHz and WISE W2 (), to analyze the distribution of star-forming galaxies and AGN in our sample. This ratio, discussed in this work for the first time, is particularly convenient for identifying radio-loud AGN, including optically quiescent radio galaxies with past jet activity, because the low-frequency radio flux is sensitive to older and less energetic synchrotron emission. On the other hand, the IR flux probes the presence of hot dust associated with the active nucleus or star formation. In star-forming galaxies this ratio would be dominated by the star-formation activity and the associated contribution from supernova remnants, while in radio-quiet QSOs low ratios are caused by the strong IR dust emission. Radio-loud sources have an extra contribution to the radio fluxes caused by the jet activity, which, at low frequencies typically comes from extended radio structures. This extra contribution distinguishes the galaxies with high radio-loudness ratios (hereafter radio-loud galaxies) from star-forming galaxies and radio-quiet QSOs. Note that although the rest-frame frequencies of this ratio are different for each redshift, our range is not large enough to avoid tracing the synchrotron continuum in radio and the dust emission in the IR. Most galaxies and a large fraction of QSOs are below . In this case, the WISE W2 band rest-frequency would correspond to , which is still representative of the AGN hot dust (e.g. Elvis et al. 1994). On the other hand, the LoTSS 144 MHz frequency would correspond to approximately 500 MHz, which still probes the synchrotron continuum. Although redshift may introduce a scatter in the radio-to-IR ratios, the latter still provides a meaningful radio-quiet and radio-loud source separation. Additionally, variability may affect the low frequency and mid- to near IR continuum differently, owing to the distinct timescales associated with the jet and hot dust components. A larger variability is expected in the latter due to the faster response to luminosity changes in the accretion disk. Nevertheless, we only expect a minor fraction of IR-variable sources in our sample (10%), as suggested by Sánchez et al. (2017) for a sample of AGN at z 2. Additionally, the radio-loudness distribution covers a range of 7 orders of magnitude, thus significant deviations are only expected for sources with extraordinarily large variability amplitudes.
II.9 Stellar masses and star formation rates
@@
Arizo-Borillo et al. (2025) obtain stellar masses, SFRs, rest-frame colors and luminosities from J-PLUS extended sources according to the sglc_prob_star parameter (López-Sanjuan et al., 2019) with the CIGALE SED fitting code. They adopt J-PLUS photometric redshifts, fluxes and flux errors and use Bruzual and Charlot (2003) stellar populations with Chabrier (2003) initial mass functions with different metallicities. We scale their mass estimates and SFRs (averaged over the last 100 Myrs), initially computed for sources with J-PLUS photo-z estimates up to , to our new z_best using the previously defined cosmology. Approximately % of the sources in our sample have a stellar mass estimate, with the majority (95%) being galaxies. We discard the values for QSOs, as AGN templates in CIGALE do not provide reliable stellar mass estimates.
@@
These results are compared with LoTSS-based estimates from Hardcastle et al. (2023), who follow the SED-fitting approach in Duncan (2022) using DESI Legacy Surveys broad-bands and WISE photometry, but introducing photo-z uncertainties. After scaling their masses to our z_best, both estimates are compared in the left panel of Fig. 5. LoTSS values show a positive median offset of 0.18 dex with 0.20 dex scatter for galaxies with reliable estimates and . The offset likely arises from differences in flux measurements and apertures, while the scatter reflects photometric uncertainties. Brighter galaxies ( and z 0.2) show a lower scatter than the median squared error of the J-PLUS mass estimates, meaning that the errors are slightly overestimated. Both offset and scatter are slightly larger for fainter galaxies ( and z 0.2). Overall, J-PLUS narrow-band derived masses are in good agreement with LoTSS broad-band estimates, confirming the robustness of our approach. The J-HERTz catalog includes stellar masses and SFRs at z_best from J-PLUS photometry. The right panel of Fig. 5 shows the distribution of stellar masses, with a few extreme outliers caused by large differences between the original J-PLUS photo-z and the newly obtained z_best, which mainly happens whtn the spectroscopic class in SDSS/DESI is not a galaxy and can be filtered with the incompatible_spec_class flag.
III Source properties in J-HERTz
III.1 QSOs
@@
In this section we analyze the main properties of the nearly 31k QSOs included in the catalog. These include their best redshift, radio luminosity, radio linear size and radio-loudness, whose distributions are shown in the orange histograms of Fig. 6. The QSO best-z distribution in our sample peaks around , in agreement with the higher density of active galaxies found at the cosmic noon (Madau and Dickinson, 2014), but it spans a wide range of values. The lowest value with a compatible spectroscopic class is 0.028 from SDSS (WISEA J160443.63+531517.8), although this appears to be a misidentification of H emission, which actually corresponds to [OII] emission at . The highest value of also comes from SDSS (WISEA J234219.09+270845.9), however spectroscopic redshifts above 5 are unreliable. Cases like either of these two examples (when the spec-z and photo-z of the same spectroscopic survey are very different, or the spec-z is larger than 5) are marked with an unreliable_spec_z flag and are not included in the histogram or the z_best parameter, although they are included in the zspec parameter (see Section IV.4). The radio luminosities peak at , with the majority falling between and . The few outliers with very low luminosities correspond to sources where the z_best comes from a low spec_z of an object spectrally classified as a star, and can be avoided with the incompatible_spec_class flag. The estimated linear size of the QSO radio emission for Resolved sources with an angular size larger than 6 arcsec mostly falls below 2700 kpc, with the most extended radio source reaching 4.3 Mpc and 20 arcmin (WISEA J093139.04+320400.1; Kuźmicz et al., 2018). We consider the host galaxies resolved if their major axis () is larger than the typical FWHM value of 1.2 arcsec in J-PLUS. By this criterion, approximately 20% of the QSO host galaxies detected by J-PLUS are resolved. Over 70% of these have an axis ratio , indicating a high degree of sphericity typical of elliptical galaxies, which are likely the usual hosts of QSOs. Taking only the cases with a relative error in and bands below 10%, QSOs have a median and dispersion () color of 0.20 0.12. The radio-loudness of QSOs is not described by a single Gaussian distribution, showing instead a bimodal distribution, i.e. with two local maxima centered around / and . We explore this further in Section IV.1.
III.2 Galaxies